OpenStudy (anonymous):

Is anyone willing to help me with 8 out of 50 questions? Please? Comment if you will, then I'll post one at a time.

OpenStudy (mathmale):

That's a big order, Florida girl. Why not post one at a time? You'd probably get more responses that way.

OpenStudy (anonymous):

Okay.. Thanks. Its only 8 questions I am not sure about or don't understand. Thanks though. @mathmale

OpenStudy (mathmale):

I wasn't trying to dismiss you. On the contrary, I was inviting you to post a question so that we could get started discussing it.

OpenStudy (anonymous):

oh! Here is the first 1. A system of equations is shown below. 4x = - 3y + 17 3x - 4y = - 6 What is the solution to this system of equations? (-2, -3) (3, 2) (2, 3) (-3, -2)

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

4x = - 3y + 17 3x - 4y = - 6 Which method of solving systems of linear equations is most familiar to you? substitution? addition/subtraction? graphing? We'll go with whichever method you choose.

OpenStudy (anonymous):

the substitution method is what I am most familiar with. Lets use that one. @mathmale

OpenStudy (mathmale):

Fine. Take a look at the first equation. 4x=-3y+17. Solve for x by dividing every term by 4. would you do that, please? When you're done, we'll substitute this expression for x into the second equation.

OpenStudy (anonymous):

Okay, divided every term by 4. 4x=-3y+17 x= -.75y+4.25 Is this correct? @mathmale

OpenStudy (mathmale):

Looks great! Now please substitute -.75y+4.25 for x in the second equation. Solve for y.

OpenStudy (anonymous):

3x-4y=-6 3(-.75+4.25)-4y=-6 -2.25+12.75-4y=-6 10.50-4y=-6 -4y=-16.50 y= -4.125 Is this all correct?

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Just about perfect! Congrats! But take a second look at -4y=-16.50. Divide both sides by -1. What do you get? How would you now find x? How would you check your final result?

OpenStudy (anonymous):

Okay.. I divided both sides by -1. 4x= 16.50 Then I would divide 4? Which would be x=4.125 I would fill x in for 4.125, right? @mathmale

OpenStudy (mathmale):

Actually, that'd be y=4.125, same as your original answer, except for sign.

OpenStudy (mathmale):

Florida: You're on the right track. Really. But I note that our y does not match any of the given answers, which means that somewhere along the line we have an arithmetic error. I will work through this quickly and see whether I can find it. If you're familiar with the addition/subtraction (elimination) method, why not try that in the meantime? In other words, solve the system 4x+3y=17 3x-4y=-6.

OpenStudy (mathmale):

Florida: I was able to solve the system of equations by substitution; I got y=3. Solving 4x = - 3y + 17 for x, I got x=(-3y+17)/4. I chose NOT to convert that to decimal fractions. I then substituted this expression for x into the second equation and got the following: \[3\left( \frac{ -3y+17 }{ 4 } \right)-4y=-6.\]

OpenStudy (mathmale):

Multiply every term by 4 to eliminate fractions. Solve the resulting equation for y. I'm so sorry, but I need to get off the computer pretty quickly. I encourage you to post your other questions too. If you're not in a hurry, I'd work with you again much later today or sometime tomorrow, but urge you not to wait for me. Good luck!

OpenStudy (anonymous):

okay, I asked my dad, since he just got home and he has walked me through it, my answer is (2,3)